### v1

# Generalizing Scaling Laws for Dense and Sparse Large Language Models

Abstract
Over the past few years, the size of language models has grown exponentially, as has the computational cost to train these large models. This rapid growth has motivated researchers to develop new techniques aimed at enhancing the efficiency of the training process. Despite these advancements, optimally predicting the model size or allocating optimal resources remains a challenge. Several efforts have addressed the challenge by proposing different scaling laws, but almost all of them are architecture-specific (dense or sparse). In this work we revisit existing scaling laws and propose a generalized scaling law to provide a unified framework that is applicable to both dense and sparse large language models. We evaluate and compare our proposed scaling law with existing scaling laws to demonstrate its effectiveness.
Index Terms:
LLMs, Dense models, Sparse models, Scaling laws, Sparsity
I Introduction
In recent years, transformer architectures [1] have revolutionized the deep learning approach. These architectures are now the foundation for the majority of popular large language models (LLMs). The reason for the widespread adoption of these architectures is the attention mechanism, which captures the deep contextual relationship in the text data. Because of the large size of LLM models and the large amount of data needed for training, however, the overall training of LLMs becomes time-consuming and expensive. As a result, determining the optimal allocation of resources, model size, and training data and estimating the training performance before conducting the actual training are crucial and have become an important topic in artificial intelligence research. Several approaches have been explored to study the scaling behavior of transformer models. The scaling laws allow researchers and developers to determine the optimal allocation of compute, to select optimal model size and data volume for training foundation models efficiently, and to estimate the performance prior to training. Until now, however, nearly all the approaches [2, 3, 4, 5, 6, 7, 8] have been architecture specific (dense or sparse). In this work, we generalize the existing empirical scaling laws and propose a generalized scaling law for both dense and sparse models that enables accurate performance prediction and optimal resource allocation across different architectures.
The field of natural language processing has undergone significant advancement with the development of transformer architectures [1] such as BERT [9] and GPT-3 [10]. Since ChatGPT was publicly released and made available in the fall of 2022, LLMs have gained more attention across diverse domains as a key innovation technology. Several models [11, 12, 13, 14, 15] have been developed over the past few years with traditional dense architecture. A variation of the transformer architecture where a fraction of available parameters are activated—namely, sparse transformers[16]—has become popular lately, and several models [17, 18, 19, 20, 21, 22, 23, 24] have been developed using the concept of sparsification. One common factor among all these models is that as they are evolving: their size have been increasing significantly from GPT-2 with $1.5$ billion parameters [25] to PaLM with 540 billion parameters [14], to DeepSeek-V3, a strong mixture-of-experts (MoE) language model with 671 billion parameters, 37 billion of which are activated for each token [26], which has 94.49% sparsity. With the increase of size, these models require large training datasets, eventually increasing the compute budget [2, 3] to several petaFLOP/s-days of computation to achieve optimal performance, thus making the models expensive to train.
In traditional transformer architectures, parameters of every layer are active for every token. Each token flows through the same feed-forward network; attention layers analyze the entire input sequence simultaneously, and every attention head considers every other position. These kinds of models are dense. The complex and exhaustive architecture allows the dense models to understand the rich contextual meaning of the input dataset. But it comes with high computational cost and scales up with both model size and sequence length.
One way to minimize the cost is to introduce sparsity in a model, where not all the parameters available in the model are used to process a token. Sparsity can be introduced in several ways. We focus on two methods: pruning [27] and mixture of experts [28]. Pruning-based sparsity permanently removes parameters from the model. On the other hand, MoE introduces new experts and uses only a fraction of them to process a token. (A more detailed explanation of pruning and MoE is given in Subsection II-B.)
Foundation models have gained attention across diverse domains because of their high accuracy, reasoning abilities, and generative power. Achieving these capabilities, however, requires vast amounts of training data and substantial compute resources. As a result, optimal allocation of compute, model size, and data volume has become critical to minimizing training costs.
Scaling laws—empirical relationships often expressed as power laws—describe a model’s performance based on allocated resources. These laws help determine the optimal model size and data volume for a given compute budget to achieve optimal pretraining performance. Several such empirical scaling laws based on different LLM architectures have been proposed in recent years.[2, 3, 4, 5, 29, 30, 7, 8, 31] Others are based on training precision [30], distillation [32], and quantization [29]. Several studies also have been done on vision [4], audio [33, 34], and mixed-modal language models [35]. This interest in scaling laws for different domains makes them pivotal in deep neural network research.
In this paper we focus on the scaling laws for dense and sparse LLMs. We make the following contributions:
  • We discuss the existing empirical scaling laws for dense and sparse LLMs and their limitations (Section II).
  • We propose a generalized scaling law for both dense and sparse LLMs and discusses the relationship between the proposed scaling law and the existing laws (Section III).
  • We evaluate and compare the performance of the proposed scaling law and the existing laws using some test models to demonstrate its effectiveness, and we present hyperparameter optimization results for some existing scaling laws (Section IV).
Section V summarizes this work and suggests avenues for potential future research.
II Scaling Laws for LLMs
Transformer models use a fully connected attention mechanism in which the token is addressed by the other tokens of the sequence. This process, which activates all the parameters in the model, is computationally expensive. The cost of training a large transformer model depends on several factors. Based on the number of parameters available in the architecture and the number of parameters that are being used during pretraining, transformer models can be categorized into dense and sparse transformers, which have shown different scaling behaviors. In the following, we discuss the scaling laws for dense and sparse LLMs.
II-A Dense Scaling Laws
Traditional transformer models are dense in nature. Kaplan et al. [2] studied the behavior of LLMs with a series of experiments and different hyperparameters: model size, model shape, total compute, data volume, and batch size. The authors observed that performance or model loss during pretraining follows a power law relationship with the total number of parameters, the total number of training tokens, and the total compute budget. They established that model loss depends strongly on scale, which consists of the number of model parameters $N$, the dataset size (number of tokens) $D$, and the amount of compute $C$ used for training. It also depends mildly on other architectural hyperparameters such as model shape: depth and width. For dense models, the compute $C$ (training FLOPs) required to train a transformer model with $N$ parameters and $D$ tokens is $C=6ND$ [2] because approximately 6 floating point operations are needed for every parameter in the model and every training sample. They suggested that if compute is increased by $10$ times, then model size should be increased by $5.5$ times and dataset size by $1.8$ times. They represented training loss as a function of number of parameters and data size and proposed the following empirical scaling law:
$L(N,D)=\left[\left(\frac{N_{C}}{N}\right)^{\frac{\alpha_{N}}{\alpha_{D}}}+\frac{D_{C}}{D}\right]^{\alpha_{D}}.$ (1)
The best-fitted values for the coefficients and exponents in Equation 1 are listed in Table I.
Coefficients $\alpha_{N}$ $\alpha_{D}$ $N_{C}$ $D_{C}$
Values 0.076 0.103 $6.4\times 10^{13}$ $1.8\times 10^{13}$
TABLE I: Fitted parameter values for Equation 1
Using three different approaches—fixed model size with varying numbers of training tokens, IsoFLOP profiling, and parametric function fitting—Hoffmann et al.  [3] extended the work in [2] and concluded that model size and total number of tokens should be increased in proportion. In other words, if the total number of parameters or model size is doubled, data volume or total number of tokens should be doubled as well. They represented training loss as a combination of irreducible loss or entropy $e$, a parameter-dependent term or the loss incurred by a model of fixed parameter size, and a data size term or the loss induced by a fixed number of training steps. Based on their observations, the authors proposed the following scaling law:
$L(N,D)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}$ (2)
This law establishes a power-law relationship between loss $L$, number of parameters $N$, and the data size $D$. Here $e$ captures the entropy of the natural text, which is the minimum loss as both $N$ and $D$ approach infinity. The second term (parameter-dependent term) captures the inverse relationship between model size $N$ and loss. The third term (data-dependent term) captures the impact of the data size $D$ on loss. The constants $a$ and $b$ and the exponents $\alpha$ and $\beta$ are determined empirically through some experiments and fitting the data. Its best-fitted values for the coefficients and exponents are listed in Table II. Equation 2 currently is a well-known empirical scaling law for dense LLMs.
Coefficients $e$ $a$ $b$ $\alpha$ $\beta$
Values $1.69$ $406.4$ $410.7$ $0.34$ $0.28$
TABLE II: Fitted parameter values for Equation 2
II-B Sparse Scaling Laws
Sparse models use only a subset of the parameters available in the models. As a result, Equation 2 needs to be adjusted and extended. Several works had proposed scaling laws for sparse transformer models [4, 5, 8]. These scaling laws are based on different sparsification methods. In this work we focus on two types of sparsifications: pruning and mixture-of-experts (MoE).
II-B1 Pruned Transformer Models
One way to introduce sparsity in transformer models is pruning, where a number of parameters are removed from the architecture and are not used during the forward pass or backpropagation, hence making no contribution to the training process. Pruning can be categorized into three main groups [27]. In structured pruning, entire higher-level units or even whole layers are removed; this process delivers universal speedup but sometimes at the cost of accuracy. In unstructured pruning, individual weights are removed from anywhere in the network; usually the least important weights are pruned, with the importance based on certain criteria. In semi-structured pruning, one tries to strike a balance between unstructured and structured pruning approaches by enforcing fixed pruning patterns in small regions of the network.
Frantar et al. [4] studied the scaling behavior of models pruned by unstructured pruning. Sparsity $S$ is defined as $S=\frac{\text{Total Parameters - Nonzero Parameters}}{\text{Total Parameters}}$. The experimental results with a total of $48$ models trained showed that sparsity introduced by pruning affects only the parameter-dependent term of Equation 2 and has no impact on the data-dependent term. The loss was fitted as a function of nonzero parameters $N$, data $D$, and sparsity $S$ as shown in the following equation.
$L(N,D,S)=\left(a_{S}(1-S)^{b_{S}}+c_{S}\right).\left(\frac{1}{N}\right)^{b_{N}}+\left(\frac{a_{D}}{D}\right)^{b_{D}}+c$ (3)
Following the similar approach of [3] to find the optimal values for the coefficients of Equation 3, the best-fitted values for these coefficients are determined empirically through some experiments and fitting the data shown in Table III.
Coefficients $a_{S}$ $b_{S}$ $c_{S}$ $b_{N}$ $a_{D}$ $b_{D}$ $c$
Values $16.8$ $0.722$ $45$ $0.245$ $6.90\times 10^{8}$ $0.203$ $0.651$
TABLE III: Fitted parameter values for Equation 3
To draw the similarities between Equations 2 and 3, we used the similar symbols in Equation 2 to reformat the coefficients in Equation 3. We used $e$ instead of $c$ to capture the lower bound on the loss that represents the inherent stochasticity of the modeling problem. We replaced $b_{N}$ and $b_{D}$ with $\alpha$ and $\beta$, respectively, and ${a_{D}}^{b_{D}}$ with $b$. The result is the following equation:
$L(N,D,S)=e+\left(a_{S}(1-S)^{b_{S}}+c_{S}\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}.$ (4)
Equation 4 helps visualize the transition from the dense scaling law Equation 2 to the sparse scaling law Equation 3 in order to investigate which factor is directly impacted by sparsity. Based on the changes in the coefficient representation, the best-fitted values need to be updated correspondingly. The updated values for the coefficients in Equation 4 are calculated and listed in Table IV.
Coefficients $a_{S}$ $b_{S}$ $c_{S}$ $b$ $\alpha$ $\beta$ $e$
Values $16.8$ $0.722$ $45$ $62.271$ $0.245$ $0.203$ $0.651$
TABLE IV: Fitted parameter values for Equation 4
While Equation 3 functionally captures the effect of sparsity on loss, the maximum sparsity used in the experiments was $87.5\%$ [4]. We have generated an estimated IsoFLOP dataset for the compute budget $1e20$ from the experiments conducted in [5], which included sparsities of more than $87.5\%$. When we used the new dataset to plot the scaling behavior of Equation 3, we noticed a spike for smaller nonzero parameters, as shown in Figure 1, where each line connects models with different sparsity and model sizes trained under the same amount of compute budget $1e20$ with nonzero parameters as x-axis and training loss as y-axis. Increasing the nonzero parameters results in decreasing the data size. However, based on the work in [4, 5], sparsity should provide a smooth performance gain as we increase sparsity. This indicates that Equation 3 does not capture the scaling behavior of highly sparse models very well.
Figure 1: IsoFLOP plot of Equation 3 for different sparsities
On the other hand, the largest model used in [4] had only $85M$ nonzero parameters. Compared with the current work, these models are relatively much smaller. This work used the scaling law Equation 2 [3] as the base for the approach. For $0\%$ sparsity, or for dense models, the loss should be equal to Equation 2. But for the same data set with no sparsity, Equation 3 provides different coefficient values from those of Equation 2, as shown in Figure 2, there is a difference of up to 2.5 in loss. Thus, we can conclude that Equation 2 is not a special case of Equation 3 if the sparsity is 0.
Figure 2: Comparison between Equations 2 and 3 for dense models
II-B2 MoE Models
Another way to introduce sparsity is to use the mixture-of-experts method [28] where the feed-forward network for a transformer model is replicated multiple times and each of these copies is called an “expert.” Only a fraction of experts are activated per token, and this number is a fixed value. The tokens are routed to different experts during the pretraining process. The size of the experts is usually equal to the size of the feed-forward network.
Abnar et al. [5] investigated the sparse MoEs through a large-scale empirical study. The sparsity $S$ was defined as $S=\frac{E-K}{E}$, where $E$ is the total number of experts and $K$ is the number of active experts. For five different compute budgets and seven different sparsities, different models were trained. Based on the experimental results, the authors proposed the following scaling law:
$L(N,D,S)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}+\frac{c}{(1-S)^{\lambda}}+\frac{d}{(1-S)^{\delta}N^{\gamma}}.$ (5)
Through a grid initialization, for a sparsity of $98\%$, the best-fitted coefficient values in Equation 5 are determined empirically through the experiments and fitting the data shown in Table V.
Coefficients $e$ $a$ $b$ $c$ $d$
Values $0.94$ $16612.50$ $5455.67$ $0.4598$ $17.26$
Coefficients $\alpha$ $\beta$ $\lambda$ $\delta$ $\gamma$
Values $0.5962$ $0.3954$ 0.1666-0.1666 $0.1603$ $0.1595$
TABLE V: Fitted parameter values for Equation 4
Equation 5 introduced two different terms in the scaling law along with the previous three terms from [3] to capture the impact of sparsity. The newly introduced terms are interdependent. The fifth term in Equation 5 can be interpreted as an active parameter term because the denominator term is a multiplication of $N$ and 1S1-S. Because the total number of parameters already includes the active parameters, however, having two different terms for total parameters and active parameters in Equation 5 causes some redundancy in the loss function.
This work also used Equation 2 as the base. But for $0\%$ sparsity we apply the models from [5] to Equations 5 and 2 to generate Figure 3 and notice a difference in loss values. This indicates that Equation 2 is not a special case of Equation 5 if the sparsity is 0.
Figure 3: Comparison between Equations 2 and 5 for dense models
III Generalizing Scaling Laws
We observed that the previous work and their scaling laws are based on the number of nonzero or active parameters. Dense models use all the parameters, and hence in [3] $N$ was used as the total number of parameters in its scaling law. In [4] the nonzero parameters $N$ were considered in its scaling law 3. In [5] the authors modified the compute budget $C=6ND$ from [2] to $C=6N_{a}D$, where $N_{a}$ is the number of active parameters. These scaling laws do not cover the dense scaling law when the sparsity is 0.
These observations and limitations motivate us to generalize the scaling laws of [3, 4, 5] into a single representation that can effectively capture the experimental behavior and features observed in these works. We want to make sure that the new scaling law works for both dense and sparse models. As we increase model size, data size, and compute, we want to get better performance [2, 3, 4, 5]. Similarly, as we increase the sparsity, we want to obtain performance improvement [4, 5]. Furthermore, for $0\%$ sparsity or dense models, we need to make sure that our scaling law behaves the same as the original scaling law in Equation 2.
For simplicity, we use the term active parameters $N$ for active/nonzero parameters in the MoE and pruned models. For the dense models, the total number of active parameters is the total number of parameters. We use Equation 2 as the base to extend it to a generalized scaling law. We define the sparsity as follows:
$S=\frac{\text{Total Parameters - Active Parameters}}{\text{Total Parameters}}$ (6)
Where $0\leq S<1$ because the active parameters in a model cannot be zero. Therefore $\frac{N}{1-S}$ is the total number of parameters. In the parameter-dependent second term in Equation 2, we replace $N$ with $\frac{N}{1-S}$; the parameter-dependent term thus becomes $\frac{a}{\left({\frac{N}{1-S}}\right)^{\alpha}}$, which is simplified as $\frac{a(1-S)^{\alpha}}{N^{\alpha}}$. Similar to the approach in [4] we introduce an upper-bound constant sparsity factor $c$, which is caused by the sparsity; but this factor, which contributes to the loss, depends on the sparsity $S$. So we define the parameter-dependent term as $\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}$. For sparsity $S=0$ or dense models, this term is the same as the original parameter-dependent term in Equation 2.
We have noticed that different approaches adjusted the entropy $e$ in different ways, as shown in Tables II, IV, and V. The value of $e$ is impacted by the sparsity. We adjust the value of $e$ by multiplying it with $(1-S)^{\gamma}$. Accumulating these changes, we propose the following scaling law:
$L(N,D,S)=e(1-S)^{\gamma}+\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}.$ (7)
This equation quantifies the impact of sparsity on loss. Notice that we did not change the values of the data-dependent term in Equation 2, because the work in [4, 5] has shown experimentally that sparsity has no effect on this term.
The fitted coefficient values in our proposed Equation 7 are listed in Table VI. Since we have used the empirical scaling law in [3] as the base, we simply apply the coefficient values in Table II to Equation 7. For the sparsity factor $c$, we use the dataset for the largest model with the largest sparsity in [4] to fit Equation 7 to obtain the value for the sparsity factor $c$. The value of $\gamma$ was empirically obtained.
Coefficients $e$ $a$ $b$ $c$ $\alpha$ $\beta$ $\gamma$
Values $1.69$ $406.4$ $410.7$ $93.45$ $0.34$ $0.28$ $10^{-2}$
TABLE VI: Fitted parameter values for law 7
Notice that Equation 7 is similar to Equation 4; however, the coefficients and the parameter-dependent term are different. If the sparsity is 0, Equation 4 is not the same as Equation 2, as shown in Figure 2.
Lemma: If the sparsity $S$ is 0, Equation 7 is the same as Equation 2.
If the sparsity $S$ is 0, the total number of parameters is equal to the total number of active parameters. From Equation 7 we have
$L(N,D,0)=e(1-0)^{\gamma}+\left(a(1-0)^{\alpha}+c\cdot 0\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}.$ (8)
Then, simplifying Equation 8, we have
$L(N,D,0)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}.$
From Equation 2, we have
$L(N,D)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}.$
Because the coefficients $e,a,b,\alpha,\beta$ are the same, we have
L(N,D,0)L(N,D)=0L(N,D,0)-L(N,D)=0.
IV Performance Results and Discussion
IV-A Performance Results
In this section we compare the performance between our proposed scaling law and the previous scaling laws for dense and sparse models. We have collected the experimental datasets (number of parameters, number of tokens, and sparsity) used in [3, 4, 5]. Hoffman et al. [3] provided a list of nine models with number of parameters and training tokens used in their experiments. Franta et al.  [4] provided four different numbers of nonzero parameters, three different numbers of training tokens, and four different sparsity levels with a total of $48$ experimental runs to collect their experimental results. Abnar et al.  [5] used five different numbers of compute budgets, seven different sparsity levels, and a range of model sizes. Using the formula $C=6ND$, we calculated their number of training tokens and listed the ranges in Table VII. We use these experimental datasets to generate plots for their respective scaling laws, and we use the same datasets to generate plots using our proposed scaling law in order to do a head-to-head comparison.
Scaling Law Parameters Tokens Sparsities(%) Test Models
Hoffman et al. [3] 400M10T400M{-}10T 8B216.2B8B{-}216.2B - 99
Frantar et al. [4] 1.3M85M1.3M{-}85M 16B65B16B{-}65B $0,50,75,87.5$ $48$
Abnar et al. [5] 329M21.2B329M{-}21.2B 15B128B15B{-}128B $0,25,50,75,90,95,98$ $35$
TABLE VII: Parameters, tokens, sparsity, and number of test models collected from the corresponding works.
When we compare our proposed scaling law in Equation 7 with the dense scaling law [3] in Equation 2, as we discussed in the lemma in the preceding section, if the sparsity is 0, Equation 7 is the same as Equation 2. Figure 4 shows the loss vs compute budget using the dataset in [3]. On the $x$-axis is compute. which depends on parameters $N$ and tokens $D$ and hence captures the impact of both ($C=6ND$). On the $y$-axis is the calculated loss. We see an identical plot, indicating that our proposed scaling law behaves the same as the original scaling law for dense models.
Figure 4: Comparison between Hoffman’s scaling law in Equation 2 and the proposed scaling law in Equation 7
We compare the proposed scaling law in Equation 7 with the scaling law [4] in Equation 3. We use the 48 models to conduct the experiments in [4] for performance comparison. Figures 5 (a) and 5 (b) show that our proposed scaling law replicates the behavior of the scaling law in [4]. We observe that with the increase of the sparsity the loss decreases. However, there is a difference between the loss values for both scaling laws which is similar to Figure 2. This is because the models used for the experiments in [4] are relatively much smaller (the maximum number of parameters just 85M) than other works, and the adjusted coefficient values in the proposed scaling law are for larger models, as shown in Figure 4 where the loss decreases with increasing the model size and compute budget. Hence we see a relatively larger loss for the much smaller datasets.
(a) Scaling behavior for Equation 3 with varying compute budget
(b) Scaling behavior for Equation 7 with varying compute budget
Figure 5: Comparison between (a) the scaling law in Equation 3 and (b) the proposed scaling law in Equation 7
We also compare our proposed scaling law in Equation 7 with the scaling law [5] in Equation 5. For performance comparison we use the interpolated values of parameters, tokens, and the exact compute budget of the 35 models the authors used to conduct the experiments in [5]. We have drawn compute budget vs loss plots in Figures 6(a) and 6(b) to compare Equations 5 and 7. We observe that with the increase of the sparsity and compute, the loss decreases; and both figures show a similar pattern. Increasing the compute budget means an increase in the active parameter size ($N_{a}$) and number of tokens ($D$). This indicates that the proposed scaling law shows scaling behavior similar to that in [5].
(a) Scaling behavior for Equation 5 with varying compute budget
(b) Scaling behavior for Equation 7 with varying compute budget
Figure 6: Comparison between (a) the scaling law in Equation 5 and (b) the proposed scaling law in Equation 7
IV-B Hyperparameter Optimization for Better Coefficients
Most of the previous scaling laws discussed in this paper used grid search to find the best values for the coefficients. While grid search is a commonly used approach, it is still inefficient. As an alternative, we have used ytopt [36, 37, 38] for hyperparameter optimization. ytopt is a Bayesian optimization–based autotuner that builds a surrogate model to explore the promising regions to find the optimal values with fewer evaluations. In our previous work [39], ytopt outperformed grid search, random search, a genetic algorithm, and XGBoost in accuracy and tuning time. Unlike traditional autotuners, ytopt leverages libEnsemble’s [40] asynchronous manager-worker framework to run multiple evaluations in parallel, reducing the one-by-one autotuning time.
To use ytopt for the optimization process for finding the best coefficients for the previous scaling laws, we use the best coefficients for the scaling laws in Tables II, IV, and V as the baseline and the default values to define the proper search space for each coefficient. We carefully define the range for each coefficient for Equations 2, 3, and 5 and then use ytopt to run the optimization process. The generated hyperparameter optimization results in Figure 7(a), (b), and (c) illustrate that ytopt identifies the best coefficient configuration (combination of coefficients) over time (blue) with the baseline (red) for the default values and that using ytopt results in lower loss values than the coefficient-based loss values (red, baseline) using grid search.
(a) Tuning Loss for Equation 2
(b) Tuning Loss for Equation 3
(c) Tuning Loss for Equation 5
Figure 7: Loss calculated by the coefficients generated by ytopt and grid search. The red lines are the loss using the best coefficients provided in the respective work, and the blue dots are the loss tuned for different evaluations over time using ytopt.
IV-C Discussion
In summary, we observe that our proposed scaling law Equation 7 effectively captures the scaling behavior of Equations 2, 3, and 5 for dense and sparse LLMs. In Figure 8 the proposed scaling law shows no spike for higher sparsity with the fixed compute budget 1e20 as we observed for Equation 3 in Figure 1. This indicates that the proposed scaling law captures the scaling behaviors for higher sparsity. For a given compute budget, the proposed scaling law can be used to estimate the best model hyperparameters with the given sparsity or to identify the optimal sparsity for the given model hyperparameters by using ytopt [36, 37].
In Figure 5 we observe that the range of loss values is different for Equations 3 and 7. This difference is mainly because the experiments done for Equation 3 use much smaller model sizes with at most 85M. For the scaling law in Equation 5 we can see that the results are almost identical. These results further confirm that our proposed scaling law captures the scaling behaviors of the MoE models.
Figure 8: IsoFLOP plot of the proposed scaling law in Equation 7 for different sparsities
In this work we have generalized the scaling behaviors of dense and sparse models. Our generalized scaling law extends and covers several existing scaling laws. Its limitations are related to the sparsity factor $c$ and the coefficient $\gamma$, which need to be further optimized for larger models. There are other ways [41, 42] to introduce sparsification in a model. Different training approaches [30, 29] may yield different scaling behavior as well. Further work can be done to incorporate these scaling laws into our generalized scaling law.
V Conclusions
Current large language models result in higher accuracies and exhibit few-shot learning and even human-like emergent abilities on a wide range of language tasks. However, these models are costly to train. As LLMs get larger and more expensive, it is crucial for researchers and developers to estimate the expected performance beforehand. Scaling laws thus are critical. But while architecture-specific scaling laws give a fine-grained estimation, as more and more new architecture and training methods are being developed, using different scaling laws for each variant becomes impractical. Our generalized scaling law covers densely activated, pruned, and MoE models. Empirical results show that our generalized approach can effectively capture the scaling behavior of the architectures. For a given compute budget, the proposed scaling law can be used to estimate the best model hyperparameters with a given sparsity or to identify the optimal sparsity for the given model hyperparameters by using ytopt.
Because LLMs inference can be memory-bound and compute-intensive, future work will extend the generalized scaling law in training to not only incorporate other scaling laws [30, 32, 29] but also take into account inference [43], especially recent agentic AI workloads for answering complex questions through reasoning using chain-of-thought prompting. Since the behavior of scaling laws varies based on training and inference approaches, these approaches can be incorporated in the generalized scaling law as well. Unifying different scaling laws into one single representation lays the foundation of architecture-agnostic representation of the scaling behavior of LLMs. As LLM models continue to diversify, this generalized approach will allow us to anticipate the return based on their allocated compute resources.
Acknowledgments
This work was supported DOE ASCR RAPIDS2 and OASIS. This material is based upon work supported by the U.S. Department of Energy, Office of Science, under contract number DE-AC02-06CH11357.